OpenStudy (anonymous):

What is the equation of the ellipse with foci at (5, 0), (-5, 0) and vertices (9, 0), (-9, 0)? Can you help explain it to me as well? :S

OpenStudy (anonymous):

Okay, do you know about the general equations of an ellipse?

OpenStudy (mertsj):

The vertices are on the x axis so the major axis is horizontal and a = 9. The focal points are (c,0) and (-c,0)

OpenStudy (anonymous):

yes .

OpenStudy (mertsj):

So c = 5. We know that a^2-b^2=c^2 or 81-b^x=25

OpenStudy (mertsj):

So b^2=56 and b=2sqrt14

OpenStudy (mertsj):

\[\frac{x^2}{81}+\frac{y^2}{56}=1\]

OpenStudy (anonymous):

Thanks I understand it better :) What is the equation of the ellipse with foci (0, 4), (0, -4) and vertices (0, 8), (0, -8)? so for this one it would be x^2/16 + y^2/64?

OpenStudy (anonymous):

x^2/16 + y^2/64 = 1

OpenStudy (mertsj):

I got 48 for b^2. How did you get 16?

OpenStudy (mertsj):

a=8, c=4 a^2-b^2=c^2 8^2-b^2=4^2

OpenStudy (mertsj):

Solve that for b^2

OpenStudy (anonymous):

64 - b^2 = 16

OpenStudy (mertsj):

yes. Finish it

OpenStudy (anonymous):

it would be 48.

OpenStudy (mertsj):

yes.

OpenStudy (anonymous):

so the answer is x^2/48 + y^2/16?

OpenStudy (mertsj):

And: \[\frac{x^2}{48}+\frac{y^2}{64}=1\]

OpenStudy (anonymous):

I have one last one..can you check if its correct or not?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

What is the equation of the ellipse with co-vertices (-20, 0), (20, 0) and foci (0, -8), (0, 8)? -20^2 + b^2 = 0 ?

OpenStudy (mertsj):

Co-vertices are (-20,0) and (20,0) so that means that b = 20 Focal points are (0,8) and (0,-8) so that means that c = 8

OpenStudy (mertsj):

a^2-20^2=8^2

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

a^2 - 400 = 64

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

so the answer is x^2/400 + y^2/64 = 1?

OpenStudy (mertsj):

\[\frac{x^2}{400}+\frac{y^2}{464}=1\]

OpenStudy (mertsj):

a^2=464

OpenStudy (anonymous):

Oh okay i get it.

OpenStudy (mertsj):

Good for you.

OpenStudy (anonymous):

Thanks for everything :)

OpenStudy (mertsj):

yw